# A countable-support symmetric iteration separating $\mathsf{PP}$ from $\mathsf{AC}$

### v4

@@ 1 Introduction
All forcing and symmetry constructions are carried out over a fixed ground universe $V$. Set-sized stages are treated in the usual set-theoretic background. To define and analyze the $\mathrm{Ord}$-length iteration as a definable class forcing (and to use the forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$), we work in a background class theory supporting the relevant recursions, namely $\mathsf{GBC}+\mathsf{ETR}$, using a class well-order of $V$ only where it is explicitly invoked for bookkeeping; see Section 2 and Remark 5.59.
@@
Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes choice for well-orderable sets: every well-orderable set of nonempty sets admits a choice function. Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds. This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used when applying Ryan–Smith localization (Theorem 4.5).
@@ 2.3 Imported iteration results from [5]
    Limit filters. In the relevant cofinality cases, the canonical limit filter generated from earlier-stage filters is normal and $\omega_{1}$-complete; see [5, §Limit filters].
@@
    Bounded-stage $\mathsf{DC}$. For set ordinals $\lambda$, assuming $V\models\mathsf{ZFC}$, if the stage filter is $\omega_{1}$-complete then the stage symmetric model $M_{\lambda}=V(\mathbb{P}_{\lambda})_{\operatorname{HS}}$ satisfies $\mathsf{DC}$; see [5, §Dependent Choice and Choice Fragments].
@@ 5 Iterating the packages over $\mathcal{N}$
    Black-box symmetric iteration. Subsection 5.5 invokes the symmetric-iteration framework to assemble the class-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and successor stages $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit package factors.
@@ 5.2 Iteration-level diagonal automorphisms
This is well-defined by Remark 5.22. In particular, for $\beta<\lambda$, $\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}$.
@@
    $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$.
@@
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection. Consequently, if $D$ is countable then so is $k\cdot D$.
@@
For $d\in D_{1}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{1})\leq\operatorname{Stab}(d)$. For $d\in D_{2}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{2})\leq\operatorname{Stab}(d)$. ∎
@@
For each $d\in D$, Lemma 5.9 provides a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$.
@@
For each $d\in D$, we have $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E_{d})\leq\operatorname{Stab}(d)$, so $(E^{\ast},D)$ is globally admissible.
@@
For each $\beta\leq\lambda$ let $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$. By Definition 5.28, $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ for every $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$. In particular, for each successor $\beta+1\leq\lambda$ the set
@@
Limit $\eta\leq\lambda$. Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently. By the limit-stage group definition ([5, Def. 3.1]), there is a unique direct-limit element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$. Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
@@
Assume $d$ has countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ in $M_{\beta}$, so $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ in $M_{\beta}$. Then $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
@@
From $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ we get that $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.28). Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the designated $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.50).
@@ 5.3 The modified limit filter
To verify clause (iii) of Definition 5.28, fix $d\in D$ and choose $k$ with $d\in D_{k}$. Then $\operatorname{Fix}(E)\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds because $(E_{k},D_{k})$ is globally admissible. Therefore $(E,D)$ is globally admissible. ∎
@@
Work in the background metatheory used to formalize class-length iterations (cf. [5, §(Class-length iterations)] and Remark 5.59). In this setting, the forcing theorem holds for the definable class forcing $\mathbb{P}_{\mathrm{Ord}}$, and the class system $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$ supports the usual recursion defining hereditarily symmetric class-names.
@@ 5.6 What the iteration yields
By Lemma 5.54(iv), the stage filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete. Therefore, by Remark 2.4(d), $M_{\alpha}\models\mathsf{DC}$. Applying $\mathsf{DC}$ inside $M_{\alpha}$ to $(A,R)$, we obtain a sequence $\langle a_{n}:n<\omega\rangle\in M_{\alpha}\subseteq\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$. Hence $\mathcal{M}\models\mathsf{DC}$. ∎
@@
Limit. Let $\lambda$ be limit and assume $M_{\beta}\models\mathsf{SVC}(S)$ for all $\beta<\lambda$. Let $X\in M_{\lambda}$ be nonempty, and choose $\dot{X}\in\operatorname{HS}_{\lambda}$ with $X=\dot{X}^{G_{\lambda}}$. By the countable-support nature of the iteration and the HS-support analysis ((cf. the countable-support analysis of HS-names in Section 3 and the preceding discussion of supports at limits), $\dot{X}$ depends on only countably many stages. Let $\langle\lambda_{n}:n<\omega\rangle$ be an increasing sequence of ordinals below $\lambda$ capturing those stages. For each $n$, let $X_{n}$ be the interpretation of the corresponding stage-truncation of $\dot{X}$; then $X_{n}\in M_{\lambda_{n}}$ and $X=\bigcup_{n<\omega}X_{n}$.
@@
$\mathsf{ZF}$ holds in $\mathcal{M}$ by Theorem 5.58. $\mathsf{DC}$ holds by Proposition 5.68. $\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.65 and 5.72. $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.72. Finally, $\neg\mathsf{AC}$ holds by Proposition 5.67. ∎
@@
By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
@@
(2) $KWP_{1}$ using $\mathsf{PP}$. By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$. Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$. Composing with the injection $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)\cong\mathcal{P}(\omega\cdot\eta)$ from (1), we obtain $X\hookrightarrow\mathcal{P}(\theta)$ for $\theta=\omega\cdot\eta$. Thus $KWP_{1}$ holds in $\mathcal{M}$.

### v5

@@ 1 Introduction
All forcing and symmetry constructions are carried out over a fixed ground universe $V$. Set-sized stages are treated in the usual set-theoretic background. To define and analyze the $\mathrm{Ord}$-length iteration as a definable class forcing (in particular, to carry out the relevant class recursions), we work in a background class theory supporting the relevant recursions, namely $\mathsf{GBC}+\mathsf{ETR}$, using a class well-order of $V$ only where it is explicitly invoked for bookkeeping; see Section 2 and Remark 5.60.
@@
Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes choice for well-orderable sets: every well-orderable family of nonempty sets admits a choice function (equivalently: for every family $\langle X_{i}:i\in I\rangle$ of nonempty sets indexed by a well-orderable set $I$, there is a choice function $c$ with $c(i)\in X_{i}$ for all $i\in I$).
Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds. This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used when applying Ryan–Smith localization (Theorem 4.5).
@@ 2.3 Imported iteration results from [5]
    Limit filters. The limit-stage filter used in this manuscript is the modified filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50). Its normality and $\omega_{1}$-completeness are proved here (Lemma 5.55). We only use [5] for the general iteration template and projection/coherence infrastructure leading to the limit stage.
@@
    Bounded-stage $\mathsf{DC}$. For any set ordinal $\lambda$, assuming $V\models\mathsf{ZFC}$, if the stage filter is $\omega_{1}$-complete then the stage symmetric model $M_{\lambda}=V(\mathbb{P}_{\lambda})_{\operatorname{HS}}$ satisfies $\mathsf{DC}$. This is proved in [5, Theorem (DC Preservation (set-length stages))].
@@ 5 Iterating the packages over $\mathcal{N}$
    Black-box symmetric iteration. Subsection 5.5 invokes the symmetric-iteration framework to assemble the class-length iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$, with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and successor stages $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$ where $\dot{\mathbb{S}}_{\alpha}$ is the finite product of the scheduled orbit package factors. By Lemma 5.65, the canonical $\mathbb{P}_{\alpha}$-names for these orbit-package iterands are hereditarily symmetric, so the two-step presentation is well-defined in the stage ground $M_{\alpha}$.
@@ 5.2 Iteration-level diagonal automorphisms
This is well-defined: if $\alpha<\beta<\lambda$, then by projection coherence (Remark 5.22),
$\pi^{\alpha}_{\beta}\bigl(\pi^{\beta}_{\lambda}(g)\bigr)=\pi^{\alpha}_{\lambda}(g),$
and applying $\pi^{0}_{\alpha}$ gives
$\pi^{0}_{\alpha}\bigl(\pi^{\beta}_{\lambda}(g)\bigr)=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\beta}(\pi^{\beta}_{\lambda}(g))\bigr)=\pi^{0}_{\alpha}\bigl(\pi^{\alpha}_{\lambda}(g)\bigr),$
so the value $\pi^{0}_{\alpha}(\pi^{\alpha}_{\lambda}(g))$ is independent of the chosen $\alpha$. In particular, for $\beta<\lambda$,
$\pi^{0}_{\beta}\circ\pi^{\beta}_{\lambda}=\pi^{0}_{\lambda}.$
@@
    $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ for every $d\in D$, where $\operatorname{Stab}_{\mathscr{G}}(d)$ denotes the stabilizer under the stage-0 action of $\mathscr{G}=\operatorname{Sym}(\omega_{1})$ on coordinates (Definition 5.24).
@@
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are coherent inverse-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection. Consequently, if $D$ is countable then so is $k\cdot D$.
@@
For $d\in D_{1}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{1})\leq\operatorname{Stab}_{\mathscr{G}}(d)$. For $d\in D_{2}$: $\operatorname{Fix}(E_{1}\cup E_{2})\leq\operatorname{Fix}(E_{2})\leq\operatorname{Stab}_{\mathscr{G}}(d)$. ∎
@@
For each $d\in D$, Lemma 5.9 provides a countable $E_{d}\subseteq\omega_{1}$ such that $\operatorname{Fix}(E_{d})\leq\operatorname{Stab}_{\mathscr{G}}(d)$.
@@
For each $d\in D$, we have $\operatorname{Fix}(E^{\ast})\leq\operatorname{Fix}(E_{d})\leq\operatorname{Stab}_{\mathscr{G}}(d)$, so $(E^{\ast},D)$ is globally admissible.
@@
For each $\beta\leq\lambda$ let $D_{\beta}:=D\cap\mathrm{Coords}_{\beta}$. By Definition 5.28, $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ for every $d\in D$, hence $\pi\cdot d=d$ for all $d\in D$. In particular, for each successor $\beta+1\leq\lambda$ the set
@@
Limit $\eta\leq\lambda$. Let $\eta$ be limit and assume $\langle k_{\beta}:\beta<\eta\rangle$ has been constructed coherently. By the limit-stage group definition ([5, Def. 3.1]), elements of $\mathscr{G}_{\eta}$ are exactly the coherent threads $\langle h_{\beta}\rangle_{\beta<\eta}$ with $\pi^{\gamma}_{\beta}(h_{\beta})=h_{\gamma}$ for $\gamma<\beta<\eta$. Since $\langle k_{\beta}\rangle_{\beta<\eta}$ is such a coherent thread by construction, it defines a unique element $k_{\eta}\in\mathscr{G}_{\eta}$ with $\pi^{\beta}_{\eta}(k_{\eta})=k_{\beta}$ for all $\beta<\eta$. Since $\widehat{\pi}^{\,D_{\eta}}$ has the same restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$, we have $k_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
@@
Assume $d$ has countable support witnessed by $E\in[\omega_{1}]^{\leq\omega}$ in $M_{\beta}$, so $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ in $M_{\beta}$. Then $\dot{\tau}^{*}\in\operatorname{HS}_{\beta+1}$; in particular, $\dot{\tau}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
@@
From $\operatorname{Fix}(E)\leq\operatorname{Stab}_{\mathscr{G}}(d)$ we get that $(E,\{d\})$ is globally admissible at stage $\beta+1$ (Definition 5.28). Hence $\Delta_{\beta+1}^{\uparrow}(E,\{d\})$ is one of the designated $\Delta^{\uparrow}$-generators of $\tilde{\mathscr{F}}^{*}_{\beta+1}$ (Definition 5.50).
@@ 5.3 The modified limit filter
Remark 5.51 (Relation to the CSI limit filter).
The companion paper [5] defines the limit-stage filter as the normal closure of head pullbacks. In this manuscript we work with the strengthened filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50) obtained by additionally adjoining the subgroups $\Delta_{\lambda}^{\uparrow}(E,D)$ for globally admissible pairs $(E,D)$.
All black-box iteration facts we import from [5] are used only in forms that require: (coherence of the projection system, normality, and $\omega_{1}$-completeness of the stage filters). These are verified for $\tilde{\mathscr{F}}^{*}_{\lambda}$ in Lemma 5.55 and its corollaries. In particular, the extra $\Delta^{\uparrow}$-generators are not present in the CSI framework and are introduced here solely to obtain the core admissibility lemma (Lemma 5.53) used in the $\neg\mathsf{AC}$ argument.
@@
To verify clause (iii) of Definition 5.28, fix $d\in D$ and choose $k$ with $d\in D_{k}$. Then $\operatorname{Fix}(E)\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}_{\mathscr{G}}(d)$, where the last inequality holds because $(E_{k},D_{k})$ is globally admissible. Therefore $(E,D)$ is globally admissible. ∎
@@
Work in the background metatheory used to formalize class-length iterations (cf. [5, §(Class-length iterations)] and Remark 5.60). In this setting, definable class recursion suffices to define the valuation map for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names and to carry out the usual recursion defining hereditarily symmetric class-names.
@@ 5.4 Bookkeeping: names versus interpretations
Lemma 5.63 (Bookkeeping guarantee).
Let $f\in\mathcal{M}$ be a surjection instance of one of our templates (PP or $\mathsf{AC}_{\mathsf{WO}}$). Choose $\dot{f}\in\operatorname{HS}$ with $f=\dot{f}^{G}$, and let $\beta:=\sup(I(\dot{f}))+1$ as in Lemma 5.67, so $f\in M_{\beta}$. Then there exists some $\alpha\geq\beta$ such that the bookkeeping schedule hits $(\beta,c_{\dot{f}})$ at stage $\alpha$, and hence:
  • either $f$ already splits in $M_{\alpha}$, in which case $f$ splits in $\mathcal{M}$;
  • or $f$ does not split in $M_{\alpha}$, and the successor-stage forcing at $\alpha$ uses the corresponding orbit package, so $f$ splits in $M_{\alpha+1}$ and hence in $\mathcal{M}$.
Proof.
By Definition 5.62 (unboundedness of the enumeration), there is some $\alpha\geq\beta$ with $\mathcal{B}(\alpha)=(\beta,c_{\dot{f}})$. At that stage, we run the bookkeeping test on $\dot{f}$ in $M_{\alpha}$. If it succeeds, the orbit package forcing at stage $\alpha$ adds a splitting object, which persists to all later stages. If it fails, then $f$ already has a splitting object in $M_{\alpha}$, and this also persists. In either case, $f$ splits in $\mathcal{M}$. ∎
@@ 5.5 The symmetric iteration (black box)
Lemma 5.65 (Stage iterands are hereditarily symmetric).
For every stage $\alpha$, the $\mathbb{P}_{\alpha}$-name $\dot{\mathbb{S}}_{\alpha}$ is an element of $\operatorname{HS}_{\alpha}$.
Proof.
If the bookkeeping test fails at stage $\alpha$, then $\dot{\mathbb{S}}_{\alpha}$ is the canonical name for the trivial forcing, which is fixed by every automorphism and hence lies in $\operatorname{HS}_{\alpha}$.
If the bookkeeping test succeeds, then $\dot{\mathbb{S}}_{\alpha}$ is (by construction) an orbit package forcing built from a scheduled surjection instance via the stage-0 action. For any $g\in\mathscr{G}_{\alpha}$, its stage-0 projection $\pi^{0}_{\alpha}(g)\in\mathscr{G}$ sends a surjection instance $f$ to $\pi^{0}_{\alpha}(g)\cdot f$, but preserves the orbit $[f]$. Consequently,
$g\cdot\dot{\mathbb{S}}_{\alpha}=\dot{\mathbb{S}}_{\alpha}.$
Thus $\mathscr{G}_{\alpha}\leq\operatorname{Stab}(\dot{\mathbb{S}}_{\alpha})$, so $\operatorname{Stab}(\dot{\mathbb{S}}_{\alpha})\in\tilde{\mathcal{F}}^{\ast}_{\alpha}$, and therefore $\dot{\mathbb{S}}_{\alpha}\in\operatorname{HS}_{\alpha}$. ∎
@@ 5.6 What the iteration yields
By Lemma 5.55(iv), the stage filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete. Since $\alpha$ is a set ordinal and $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete, the hypotheses of Remark 2.4(d) are satisfied, so $M_{\alpha}\models\mathsf{DC}$.
Applying $\mathsf{DC}$ inside $M_{\alpha}$ to $(A,R)$, we obtain a sequence $\langle a_{n}:n<\omega\rangle\in M_{\alpha}\subseteq\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$. Hence $\mathcal{M}\models\mathsf{DC}$. ∎
Definition 5.72 (Stage truncation of names).
Let $\lambda$ be a limit stage and let $\alpha<\lambda$. For a $\mathbb{P}_{\lambda}$-name $\dot{x}$, define the $\mathbb{P}_{\alpha}$-name $\dot{x}\!\upharpoonright\!\alpha$ recursively by
$\dot{x}\!\upharpoonright\!\alpha\;:=\;\Bigl\{\,\bigl\langle\dot{y}\!\upharpoonright\!\alpha,\ p\!\upharpoonright\!\alpha\bigr\rangle:\ \langle\dot{y},p\rangle\in\dot{x}\ \wedge\ \operatorname{supp}(p)\subseteq\alpha\,\Bigr\}.$
(Thus we keep only those pairs whose witnessing condition already lies in $\mathbb{P}_{\alpha}$; in that case $p\upharpoonright\alpha=p$.)
Lemma 5.73 (Union of truncations along an increasing stage sequence).
Let $\lambda$ be a limit stage and let $\dot{X}\in\operatorname{HS}_{\lambda}$ with $1\Vdash_{\mathbb{P}_{\lambda}}$$\dot{X}$ is a set”. Suppose $\langle\lambda_{n}:n<\omega\rangle$ is an increasing sequence of ordinals below $\lambda$ such that for every $\langle\dot{y},p\rangle\in\operatorname{tc}(\dot{X})$ there is $n<\omega$ with $\operatorname{supp}(p)\subseteq\lambda_{n}$. Then for every $V$-generic $G_{\lambda}$,
$\dot{X}^{G_{\lambda}}\;=\;\bigcup_{n<\omega}\bigl(\dot{X}\!\upharpoonright\!\lambda_{n}\bigr)^{G_{\lambda_{n}}}.$
Proof.
Fix $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ generic and write $G_{\lambda_{n}}:=G_{\lambda}\cap\mathbb{P}_{\lambda_{n}}$. For each $n<\omega$, set $X_{n}:=(\dot{X}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}$.
We first note (by a straightforward induction on rank) that if $\operatorname{supp}(p)\subseteq\lambda_{n}$ and $\dot{y}$ is any $\mathbb{P}_{\lambda}$-name, then
$(\dot{y}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}\;=\;\dot{y}^{G_{\lambda}}\quad\text{whenever }p\in G_{\lambda_{n}}.$
($\subseteq$). Let $x\in\dot{X}^{G_{\lambda}}$. Choose $\langle\dot{y},p\rangle\in\dot{X}$ with $p\in G_{\lambda}$ and $x=\dot{y}^{G_{\lambda}}$. By hypothesis, pick $n<\omega$ with $\operatorname{supp}(p)\subseteq\lambda_{n}$. Then $p\in\mathbb{P}_{\lambda_{n}}$ and hence $p\in G_{\lambda_{n}}$. By definition of stage truncation, $\langle\dot{y}\!\upharpoonright\!\lambda_{n},\ p\rangle\in\dot{X}\!\upharpoonright\!\lambda_{n}$, so
$x=\dot{y}^{G_{\lambda}}=(\dot{y}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}\in X_{n}.$
($\supseteq$). Let $x\in X_{n}$. Then for some $\langle\dot{y}^{\prime},p\rangle\in\dot{X}\!\upharpoonright\!\lambda_{n}$ with $p\in G_{\lambda_{n}}$ we have $x=(\dot{y}^{\prime})^{G_{\lambda_{n}}}$. By definition of $\dot{X}\!\upharpoonright\!\lambda_{n}$, there is $\langle\dot{y},p\rangle\in\dot{X}$ with $\operatorname{supp}(p)\subseteq\lambda_{n}$ and $\dot{y}^{\prime}=\dot{y}\!\upharpoonright\!\lambda_{n}$. Since $p\in G_{\lambda_{n}}\subseteq G_{\lambda}$, we conclude
$x=(\dot{y}\!\upharpoonright\!\lambda_{n})^{G_{\lambda_{n}}}=\dot{y}^{G_{\lambda}}\in\dot{X}^{G_{\lambda}}.$
This proves $\dot{X}^{G_{\lambda}}=\bigcup_{n<\omega}X_{n}$. ∎
@@
Limit. Let $\lambda$ be limit and assume $M_{\beta}\models\mathsf{SVC}(S)$ for all $\beta<\lambda$. Let $X\in M_{\lambda}$ be nonempty, and choose $\dot{X}\in\operatorname{HS}_{\lambda}$ with $X=\dot{X}^{G_{\lambda}}$.
By the countable-support analysis of HS-names, fix an increasing sequence $\langle\lambda_{n}:n<\omega\rangle$ of ordinals below $\lambda$ satisfying the hypothesis of Lemma 5.73. For each $n$, set
$X_{n}\;:=\;\bigl(\dot{X}\!\upharpoonright\!\lambda_{n}\bigr)^{G_{\lambda_{n}}}.$
Then $X_{n}\in M_{\lambda_{n}}$ and $X=\bigcup_{n<\omega}X_{n}$ by Lemma 5.73.
@@
$\mathsf{ZF}$ holds in $\mathcal{M}$ by Theorem 5.59. $\mathsf{DC}$ holds by Proposition 5.71. $\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.68 and 5.77. $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.77. Finally, $\neg\mathsf{AC}$ holds by Proposition 5.70. ∎
@@
By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
@@
(2) $KWP_{1}$ using $\mathsf{PP}$. By Theorem 4.5 together with Propositions 5.68 and 5.77, $\mathcal{M}\models\mathsf{PP}$. Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$. Composing with the injection $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)\cong\mathcal{P}(\omega\cdot\eta)$ from (1), we obtain $X\hookrightarrow\mathcal{P}(\theta)$ for $\theta=\omega\cdot\eta$. Thus $KWP_{1}$ holds in $\mathcal{M}$.